what is the slope and intercept of y=7/8x-5/9?
Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
y = (7/8)x - 5/9 | slope m = 7/8 and the y-intercept is Pt(0,-5/9)